ON THE PROFINITE TOPOLOGY ON DOUBLES OF GROUPS, II

RITA GITIK

Abstract. We show that quasiconvex subgroups in doubles of certain negatively curved LERF groups are closed in the profinite topology. This implies the existence of a large family of LERF hyperbolic 3-manifolds with boundary.